OpenStudy (anonymous):

solve 4^(x+7)=6^(x-1) x ≈ 0.04 x ≈ 0.05 x ≈ 19.51 x ≈ 28.35

OpenStudy (anonymous):

i got a but i think i am wrong

OpenStudy (precal):

log both sides use the log property to drop the exponents, then solve

OpenStudy (anonymous):

i...iam ashamed to say i dont know how to do this

OpenStudy (anonymous):

i think i did the whole thing wrong

OpenStudy (precal):

ok I will help you, no shame at all. give me a second to type it in

OpenStudy (anonymous):

thank u

OpenStudy (precal):

\[4^\left( x+7 \right)=6^\left( x-1 \right)\] \[log 4^\left( x+7 \right)=log 6^\left( x-1 \right)\]

OpenStudy (precal):

you can take the common log or natural log of both sides, method does not matter

OpenStudy (precal):

(x+7)log4=(x-1)log6 xlog4+7log4=xlog6-log6

OpenStudy (anonymous):

o.o log is on the calculator but its not in the problem is it just assumed?

OpenStudy (precal):

no we do not need that but you can use log in the calculator later to give us a number. We are just doing algebra at this point

OpenStudy (precal):

xlog4+7log4=xlog6-log6 xlog4-xlog6=-7log4-log6 x(log4-log6)=-7log4-log6 \[x=\frac{-7\log4-\log6}{\log4-\log6}\] ok put that in your calculator, be careful because you will need many parenthesis

OpenStudy (precal):

x is approximately 28.35218066

OpenStudy (precal):

Let me check my solution to see that it works

OpenStudy (anonymous):

ok :D

OpenStudy (precal):

yes it works........ :)

OpenStudy (precal):

Did you understand my steps?